The income of families in a slum have a normal distribution with a mean of $ 52 per year and a standard deviation of $ 5 . find (10mks) i)Percentage of families with an income of more than $50 per year ii). Percentage of families with an income of between $ 48 and $ 54 per year iii) .In a sample of 500 families in this slum, how many have an income of more than $ 50 per year?
1
Expert's answer
2020-02-24T12:23:36-0500

Let X=the income of families in a slum. "X\sim N(\mu, \sigma^2)"

Then


"Z={X-\mu \over \sigma}\sim N(0, 1)"

Given that "\mu=52, \sigma=5"

i)


"P(X>50)=1-P(X\leq50)="

"=1-P(Z\leq{50-52 \over 5})=1-P(Z\leq-0.4)\approx"

"\approx1-0.3446\approx0.6554"

"65.54\%"


ii)


"P(48<x<54)=P(X<54)-P(X\leq48)="

"=P(Z<{54-52 \over 5})-P(Z\leq{48-52 \over 5})="

"=P(Z<0.4)-P(Z\leq-0.8)\approx0.65542-0.21186\approx"

"\approx0.4436"

"44.36\%"


iii)

Given that "n=500"


"P(X>50)\approx0.6554"

"500\cdot0.6554\approx328"

Approximately 328 families in this slum have an income of more than $ 50 per year.